# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ IV-A Optimizing BD-RIS: A Simple SISO Case
Consider a BD-RIS-aided SISO system, consisting of a single-antenna transmitter, a single-antenna receiver, and an $M$-element BD-RIS. Define the channels from transmitter to receiver, from the transmitter to BD-RIS, and from BD-RIS to receiver, respectively, as $h_{RT}\in\mathbb{C}$, $\mathbf{h}_{IT}\in\mathbb{C}^{M\times 1}$, and $\mathbf{h}_{RI}\in\mathbb{C}^{1\times M}$, which are perfectly known at the transmitter. According to [35], assuming perfect matching, no mutual coupling, no specular reflection, and unilateral approximation at all devices, the overall channel is given by $h(\mathbf{\Theta})=h_{RT}+\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{h}_{IT}$.
@@ IV-A.1 Solutions for Unitary $\mathbf{\Theta}$
  • Closed-Form Solution: As explained in Section II, given two unitary matrices $\mathbf{V}_{RI}\in\mathbb{C}^{M\times M}$ and $\mathbf{U}_{IT}\in\mathbb{C}^{M\times M}$ with $[\mathbf{V}_{RI}]_{:,1}=\frac{\mathbf{h}_{RI}^{\mathsf{H}}}{\|\mathbf{h}_{RI}\|_{2}}$ and $[\mathbf{U}_{IT}]_{:,1}=\frac{\mathbf{h}_{IT}}{\|\mathbf{h}_{IT}\|_{2}}$, it is easy to have a global optimal solution of problem (44) as
@@
  • Orthogonal Rotation: Following the idea in [48], any unitary matrix $\mathbf{\Theta}$ can be expressed using the following parameterization
@@
    $[\mathbf{R}_{m,n}]_{i,j}=\begin{cases}1,&i=j,i\neq m,n,\\ \cos\phi_{m,n},&i=j,i=m,n,\\ \sin\phi_{m,n}e^{\jmath\psi_{m,n}},&i=m,j=n,\\ -\sin\phi_{m,n}e^{-\jmath\psi_{m,n}},&i=n,j=m,\\ 0,&\text{otherwise}.\end{cases}$ (47)
    As such, the original problem is transformed into an unconstrained optimization containing $M(M-1)$ rotation parameters of the corresponding rotation matrices, which can be directly solved by some searching-based methods, e.g., quasi-Newton methods.
  • Manifold: It is worth noting that the constraint $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$ essentially forms a Manifold [49]. As such, a typical solution is to adopt the Manifold theory to construct an unconstrained optimization on the Manifold space, such that searching-based methods, such as Gradient-Descent, can be used on the Manifold space [50].
@@ IV-A.2 Solutions for Unitary and Symmetric $\mathbf{\Theta}$
  • Matrix Decomposition: One way to decouple the constraints of $\mathbf{\Theta}$ is to decompose $\mathbf{\Theta}$ as $\mathbf{\Theta}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ [51, 52], where $\mathbf{U}\in\mathbb{R}^{M\times M}$ is orthonormal, that is $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{I}_{M}$ and $\mathbf{D}\in\mathbb{C}^{M\times M}$ is a diagonal matrix whose diagonal entries have unit modulus, that is $\mathbf{D}=\mathsf{diag}(D_{1},\ldots,D_{M})$, $|D_{m}|=1$, $\forall m\in\mathcal{M}$. Then the design of $\mathbf{\Theta}$ is transformed to the design of $\mathbf{D}$ and $\mathbf{U}$:
@@
    both having closed-form solutions as detailed in [51, 52]. Alternatively, $\mathbf{\Theta}$ can be decomposed as $\mathbf{\Theta}=\mathbf{\Psi}\mathbf{\Psi}^{\mathsf{T}}$ [53]. As such, the design of a unitary and symmetric matrix $\mathbf{\Theta}$ is equivalent to that of a unitary matrix $\mathbf{\Psi}\in\mathbb{C}^{M\times M}$, that is
@@
    which can be obtained in closed form based on the knowledge of channels as detailed in [53], or by directly using the Manifold method.
@@
    and then applies the symmetric unitary projection to satisfy the original constraints [54]. That is, with $\mathbf{\Theta}_{\mathsf{opt}}$ as the solution to (51) as detailed in [54], its symmetric projection is given by
@@
    Then, $\mathbf{\Theta}_{\mathsf{sym}}$ is further projected to the unitary domain by solving an orthogonal Procrustes problem [55], i.e.,
@@
    where $\mathbf{U}_{1}\in\mathbb{C}^{M\times M}$ and $\mathbf{U}_{2}\in\mathbb{C}^{M\times M}$ are unitary matrices from the singular value decomposition (SVD) of $\mathbf{\Theta}_{\mathsf{sym}}$, i.e., $\mathbf{\Theta}_{\mathsf{sym}}=\mathbf{U}_{1}\mathbf{\Sigma}\mathbf{U}_{2}^{\mathsf{H}}$ with $\mathbf{\Sigma}$ being a diagonal matrix containing the singular values of $\mathbf{\Theta}_{\mathsf{sym}}$ in a decreasing order.
  • Introducing Auxiliary Variables: Another direct way to decouple the two constraints of $\mathbf{\Theta}$ is to introduce an auxiliary variable $\mathbf{\Phi}=\mathbf{\Theta}$ [56], such that $\mathbf{\Theta}$ is only subject to the unitary constraint and $\mathbf{\Phi}$ is only subject to the symmetric constraint, or vise versa. In this way, the equality constraint $\mathbf{\Phi}=\mathbf{\Theta}$ can be penalized to construct the associated Lagrangian function and problem (48) can be transformed into a double-variable optimization
@@
    where $\rho_{1}>0$ is the penalty parameter and $\mathbf{\Lambda}\in\mathbb{C}^{M\times M}$ is the dual variable. Then, the two variables can be alternatively designed. Specifically, the sub-problem for $\mathbf{\Theta}$ with given $\mathbf{\Phi}$ is essentially an unconstrained optimization when focusing only on the diagonal and upper-triangular (or equivalently lower-triangular) entries of $\mathbf{\Theta}$. As such, some well-known searching-based methods for unconstrained optimization, such as Gradient-Descent, can be directly adopted. The sub-problem for $\mathbf{\Phi}$ with given $\mathbf{\Theta}$ is again an orthogonal Procrustes problem [55], whose closed-form solution can be obtained by performing SVD to $\rho_{1}\mathbf{\Theta}+\mathbf{\Lambda}$.
Remark 6: The matrix decomposition $\mathbf{\Theta}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ [51, 52] provides interesting insights to understand the relationship between BD-RIS and D-RIS. That is, a BD-RIS can be decoupled as a power divider controlled by $\mathbf{U}$ and a phase shifter network characterized by $\mathbf{D}$. In D-RIS, we have $\mathbf{U}=\mathbf{I}_{M}$, indicating that signals impinging on one element can only be reflected by the same one. However, in BD-RIS, a flexible power divider is enabled by inter-element connections such that the waves impinging on one element can (partially) flow to other elements and further be reflected.
@@ IV-A.3 Solutions for $\mathbf{Y}_{I}$
where $\mathcal{Y}$ denotes the constraint of $\mathbf{Y}_{I}$ varying according to architectures illustrated in Section III-C. The main difficulty in optimizing BD-RIS here lies in the matrix inverse due to the nonlinear mapping between $\mathbf{\Theta}$ and $\mathbf{Y}_{I}$. To tackle this difficulty, there are in general three strategies.
  • Closed-Form Solutions: As illustrated in Section II, the channel strength in (55) is upper-bounded by
    $|\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{h}_{IT}|^{2}\leq\|\mathbf{h}_{RI}\|_{2}^{2}\|\mathbf{h}_{IT}\|_{2}^{2},$ (56)
    where the equality is achieved when
@@
    It turns out that it is sufficient to solve linear equations in (58) as the global optimal solution of (55). Note that with a given architecture, one can always focus only on the nonzero entries in $\mathbf{Y}_{I}$ and solve (58). Such a closed-form solution for tree/forest-connected architectures has been derived in [40].
  • Searching-Based Methods: The optimization problem (55) is essentially unconstrained by focusing only on the upper-triangular (or equivalently lower-triangular) and diagonal entries of $\mathbf{Y}_{I}$. This can be directly solved by some searching-based methods, such as the quasi-Newton method [26].
  • Introducing Auxiliary Variables: One can introduce an auxiliary variable $\mathbf{u}=\mathbf{h}_{RI}\mathbf{\Theta}$, such that the matrix inverse can be eliminated by transferring it to a bilinear constraint [57]:
@@
    where $\rho_{2}>0$ denotes the penalty parameter and $\bm{\lambda}\in\mathbb{C}^{1\times M}$ denotes the dual variable. Then the two variables can be alternatively optimized, where the sub-problem for $\mathbf{u}$ is unconstrained and the sub-problem for $\mathbf{Y}_{I}$ is also unconstrained when focusing on optimizing only the non-zero entries in $\mathbf{Y}_{I}$.
@@
  Construct $\mathbf{\Theta}$ by two unitary matrices based on $\mathbf{h}_{RI}$ and $\mathbf{h}_{IT}$, i.e., $\mathbf{\Theta}=\mathbf{V}_{RI}\mathbf{U}_{IT}^{\mathsf{H}}$
       with $[\mathbf{V}_{RI}]_{:,1}=\frac{\mathbf{h}_{RI}^{\mathsf{H}}}{\|\mathbf{h}_{IT}\|_{2}}=\bar{\mathbf{h}}_{RI}^{\mathsf{H}}$ and $[\mathbf{U}_{IT}]_{:,1}=\frac{\mathbf{h}_{IT}}{\|\mathbf{h}_{IT}\|_{2}}=\bar{\mathbf{h}}_{IT}$.
  Decouple $\mathbf{\Theta}$ as [48] $\mathbf{\Theta}=\mathbf{\Theta}_{0}\prod_{m=1}^{M-1}\prod_{n=m+1}^{M}\mathbf{R}_{m,n}$, transform to an
       unconstrained optimization and use searching-based methods,
       e.g., quasi-Newton method.
  Adopt the Manifold theory to construct a Manifold and transform the constrained
       optimization to unconstrained optimization on Manifold space, on which
       the searching-based methods, e.g., Gradient-Descent, can be used [36, 50].
Fully/group-connected architectures with $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$ and $\mathbf{\Theta}=\mathbf{\Theta}^{\mathsf{T}}$ Coupled unitary and symmetric constraints Decouple two constraints by
  matrix decomposition $\mathbf{\Theta}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ with $\mathcal{T}^{\prime}$ [51, 52] or $\mathbf{\Theta}=\mathbf{\Psi}\mathbf{\Psi}^{\mathsf{T}}$ with $\mathcal{T}^{\prime\prime}$ [53];
  symmetric and unitary projections by $\mathbf{\Theta}_{\mathsf{sym}}=\frac{1}{2}(\mathbf{\Theta}_{\mathsf{opt}}+\mathbf{\Theta}_{\mathsf{opt}}^{\mathsf{T}})$ and solving
       orthogonal Procrustes problem $\mathbf{\Theta}=\arg\min_{\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}}~{}\|\mathbf{\Theta}-\mathbf{\Theta}_{\mathsf{sym}}\|_{\mathsf{F}}^{2}$ [54];
  introducing auxiliary variables $\mathbf{\Phi}=\mathbf{\Theta}$, with $\mathbf{\Phi}=\mathbf{\Phi}^{\mathsf{T}}$ and $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$,
       or vise versa [56].
@@
  Extract non-zero entries in $\mathbf{Y}_{I}$ and globally solve linear equations
       $\mathbf{Y}_{I}(\bar{\mathbf{h}}_{RI}^{\mathsf{H}}+\bar{\mathbf{h}}_{IT})=Y_{0}(\bar{\mathbf{h}}_{IT}-\bar{\mathbf{h}}_{RI}^{\mathsf{H}})$.
  Searching-based methods, e.g., quasi-Newton method [26], which are not
       sensitive to matrix structures.
  Introduce auxiliary variables, e.g., $\mathbf{u}=\mathbf{h}_{RI}\mathbf{\Theta}$[57], to eliminate the
       matrix inverse by $\mathbf{\Theta}=(Y_{0}\mathbf{I}_{M}+\mathbf{Y}_{I})^{-1}(Y_{0}\mathbf{I}_{M}-\mathbf{Y}_{I})\Rightarrow$
       $\mathbf{u}(Y_{0}\mathbf{I}_{M}+\mathbf{Y}_{I})=\mathbf{h}_{RI}(Y_{0}\mathbf{I}_{M}-\mathbf{Y}_{I})$.

### v2

@@ IV-A Optimizing BD-RIS: A Simple SISO Case
Consider a BD-RIS-aided SISO system, consisting of a single-antenna transmitter, a single-antenna receiver, and an $M$-element BD-RIS. Define the channels from transmitter to receiver, from the transmitter to BD-RIS, and from BD-RIS to receiver, respectively, as $h_{RT}\in\mathbb{C}$, $\mathbf{h}_{IT}\in\mathbb{C}^{M\times 1}$, and $\mathbf{h}_{RI}\in\mathbb{C}^{1\times M}$, which are perfectly known at the transmitter. According to [universal2024], assuming perfect matching, no mutual coupling, no specular reflection, and unilateral approximation at all devices, the overall channel is given by $h(\mathbf{\Theta})=h_{RT}+\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{h}_{IT}$.
@@ IV-A.1 Solutions for Unitary $\mathbf{\Theta}$
  • Closed-Form Solution: Consider the matrix $\mathbf{V}_{RI}\in\mathbb{C}^{M\times M}$ as a unitary matrix containing the right singular vectors of $\mathbf{h}_{RI}$ in its columns. In other words, the first column of $\mathbf{V}_{RI}$ is $\mathbf{V}_{RI}]_{:,1}=\frac{\mathbf{h}_{RI}^{\mathsf{H}}}{\|\mathbf{h}_{RI}\|_{2}}$ and the other columns of $\mathbf{V}_{RI}$ are unit-norm vectors mutually orthogonal and orthogonal to $\mathbf{V}_{RI}]_{:,1}$. Similarly, consider the matrix $\mathbf{U}_{IT}\in\mathbb{C}^{M\times M}$ as a unitary matrix containing the left singular vectors of $\mathbf{h}_{IT}$ in its columns. In other words, the first column of $\mathbf{U}_{IT}$ is $\mathbf{U}_{IT}]_{:,1}=\frac{\mathbf{h}_{IT}}{\|\mathbf{h}_{IT}\|_{2}}$ and the other columns of $\mathbf{U}_{IT}$ are unit-norm vectors mutually orthogonal and orthogonal to $\mathbf{U}_{IT}]_{:,1}$. Given $\mathbf{V}_{RI}$ and $\mathbf{U}_{IT}$, it is easy to see that a unitary matrix $\boldsymbol{\Theta}$ globally solving problem (42) is
@@
    The computational complexity mainly comes from the matrix multiplication and is thus $\mathcal{O}(M^{3})$.
  • Orthogonal Rotation: Following the idea in [de2010optimized], any unitary matrix $\mathbf{\Theta}$ can be expressed using the following parameterization
@@
    $\mathbf{R}_{m,n}]_{i,j}=\begin{cases}1,&i=j,i\neq m,n,\\ \cos\phi_{m,n},&i=j,i=m,n,\\ \sin\phi_{m,n}e^{\jmath\psi_{m,n}},&i=m,j=n,\\ -\sin\phi_{m,n}e^{-\jmath\psi_{m,n}},&i=n,j=m,\\ 0,&\text{otherwise}.\end{cases}$ (45)
    As such, the original problem is transformed into an unconstrained optimization containing M(M1)M(M-1) rotation parameters of the corresponding rotation matrices, which can be directly solved by some searching-based methods, e.g., quasi-Newton methods. This solution requires around $\mathcal{O}(I_{1}M^{2}(M-1)^{2})$ complexity due to the use of quasi-Newton method to solve M(M1)M(M-1) variables, where $I_{1}$ denotes the number of iterations.
  • Manifold: It is worth noting that the constraint $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$ essentially forms a Manifold [lee2018introduction]. As such, a typical solution is to adopt the Manifold theory to construct an unconstrained optimization on the Manifold space, such that searching-based methods, such as Gradient-Descent, can be used on the Manifold space [zhao2024channel]. This solution requires $\mathcal{O}(I_{2}M^{3})$ complexity, where $I_{2}$ denotes the number of iterations.
@@ IV-A.2 Solutions for Unitary and Symmetric $\mathbf{\Theta}$
  • Matrix Decomposition: One way to decouple the constraints of $\mathbf{\Theta}$ is to decompose $\mathbf{\Theta}$ as $\mathbf{\Theta}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ [nerini2023closed, sun2024new], where $\mathbf{U}\in\mathbb{R}^{M\times M}$ is orthonormal, that is $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{I}_{M}$ and $\mathbf{D}\in\mathbb{C}^{M\times M}$ is a diagonal matrix whose diagonal entries have unit modulus, that is $\mathbf{D}=\mathsf{diag}(D_{1},\ldots,D_{M})$, $|D_{m}|=1$, $\forall m\in\mathcal{M}$. Then the design of $\mathbf{\Theta}$ is transformed to the design of $\mathbf{D}$ and $\mathbf{U}$:
@@
    both having closed-form solutions as detailed in [nerini2023closed, sun2024new]. Alternatively, $\mathbf{\Theta}$ can be decomposed as $\mathbf{\Theta}=\mathbf{\Psi}\mathbf{\Psi}^{\mathsf{T}}$ [santamaria2023snr]. As such, the design of a unitary and symmetric matrix $\mathbf{\Theta}$ is equivalent to that of a unitary matrix $\mathbf{\Psi}\in\mathbb{C}^{M\times M}$, that is
@@
    which can be obtained in closed form based on the knowledge of channels as detailed in [santamaria2023snr]. The matrix decomposition requires around $\mathcal{O}(M^{3})$ due to the use of eigenvalue and Takagi’s decomposition.
@@
    and then applies the symmetric unitary projection to satisfy the original constraints [fang2023low]. That is, with $\mathbf{\Theta}_{\mathsf{opt}}$ as the solution to (49) as detailed in [fang2023low], its symmetric projection is given by
@@
    Then, $\mathbf{\Theta}_{\mathsf{sym}}$ is further projected to the unitary domain by solving an orthogonal Procrustes problem [manton2002optimization], i.e.,
@@
    where $\mathbf{U}_{1}\in\mathbb{C}^{M\times M}$ and $\mathbf{U}_{2}\in\mathbb{C}^{M\times M}$ are unitary matrices from the singular value decomposition (SVD) of $\mathbf{\Theta}_{\mathsf{sym}}$, i.e., $\mathbf{\Theta}_{\mathsf{sym}}=\mathbf{U}_{1}\mathbf{\Sigma}\mathbf{U}_{2}^{\mathsf{H}}$ with $\mathbf{\Sigma}$ being a diagonal matrix containing the singular values of $\mathbf{\Theta}_{\mathsf{sym}}$ in a decreasing order. The projection requires around $\mathcal{O}(M^{3})$ due to the use of SVD.
  • Introducing Auxiliary Variables: Another direct way to decouple the two constraints of $\mathbf{\Theta}$ is to introduce an auxiliary variable $\mathbf{\Phi}=\mathbf{\Theta}$ [zhou2023optimizing], such that $\mathbf{\Theta}$ is only subject to the unitary constraint and $\mathbf{\Phi}$ is only subject to the symmetric constraint, or vise versa. In this way, the equality constraint $\mathbf{\Phi}=\mathbf{\Theta}$ can be penalized to construct the associated Lagrangian function and problem (46) can be transformed into a double-variable optimization
@@
    where $\rho_{1}>0$ is the penalty parameter and $\mathbf{\Lambda}\in\mathbb{C}^{M\times M}$ is the dual variable. Then, the two variables can be alternatively designed. Specifically, the sub-problem for $\mathbf{\Theta}$ with given $\mathbf{\Phi}$ is essentially an unconstrained optimization when focusing only on the diagonal and upper-triangular (or equivalently lower-triangular) entries of $\mathbf{\Theta}$. As such, some well-known searching-based methods for unconstrained optimization, such as Gradient-Descent, can be directly adopted. The sub-problem for $\mathbf{\Phi}$ with given $\mathbf{\Theta}$ is again an orthogonal Procrustes problem [manton2002optimization], whose closed-form solution can be obtained by performing SVD to $\rho_{1}\mathbf{\Theta}+\mathbf{\Lambda}$. The overall complexity of such a solution depends on the specific methods used to solve the sub-problem for a symmetric $\mathbf{\Theta}$.
Remark 7: The matrix decomposition $\mathbf{\Theta}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ [nerini2023closed, sun2024new] provides interesting insights to understand the relationship between BD-RIS and D-RIS. That is, a BD-RIS can be decoupled as a power divider controlled by $\mathbf{U}$ and a phase shifter network characterized by $\mathbf{D}$. In D-RIS, we have $\mathbf{U}=\mathbf{I}_{M}$, indicating that signals impinging on one element can only be reflected by the same one. However, in BD-RIS, a flexible power divider is enabled by inter-element connections such that the waves impinging on one element can (partially) flow to other elements and further be reflected.
@@ IV-A.3 Solutions for $\mathbf{Y}_{I}$
where $\mathcal{Y}$ denotes the constraint of $\mathbf{Y}_{I}$ varying according to architectures illustrated in Section III-D. The main difficulty in optimizing BD-RIS here lies in the matrix inverse due to the nonlinear mapping between $\mathbf{\Theta}$ and $\mathbf{Y}_{I}$. To tackle this difficulty, there are in general three strategies.
  • Closed-Form Solutions: As illustrated in Section II, the channel strength in (53) is upper-bounded by $|\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{h}_{IT}|^{2}\leq\|\mathbf{h}_{RI}\|_{2}^{2}\|\mathbf{h}_{IT}\|_{2}^{2}$, where the equality is achieved when
@@
    It turns out that it is sufficient to solve linear equations in (55) as the global optimal solution of (53). Note that with a given architecture, one can always focus only on the nonzero entries in $\mathbf{Y}_{I}$ and solve (55). Such a closed-form solution for tree/forest-connected architectures has been derived in [nerini2024beyond], which requires $\mathcal{O}(M^{3})$ complexity due to the matrix inversion.
  • Searching-Based Methods: The optimization problem (53) is essentially unconstrained by focusing only on the upper-triangular (or equivalently lower-triangular) and diagonal entries of $\mathbf{Y}_{I}$. This can be directly solved by some searching-based methods, such as the quasi-Newton method [shen2021], which requires $\mathcal{O}(I_{3}\frac{M^{2}(M+1)^{2}}{4})$ to solve $\frac{M(M+1)}{2}$ variables, where $I_{3}$ denotes the number of iterations.
  • Introducing Auxiliary Variables: One can introduce $\mathbf{u}=\mathbf{h}_{RI}\mathbf{\Theta}$, such that the matrix inverse can be eliminated by transferring it to a bilinear constraint [wu2024optimization]:
@@
    where $\rho_{2}>0$ denotes the penalty parameter and $\bm{\lambda}\in\mathbb{C}^{1\times M}$ denotes the dual variable. Then the two variables can be alternatively optimized, where the sub-problem for $\mathbf{u}$ is unconstrained and the sub-problem for $\mathbf{Y}_{I}$ is also unconstrained when focusing on optimizing only the non-zero entries in $\mathbf{Y}_{I}$. The overall complexity of this solution depends on the specific methods used to solve two unconstrained sub-problems.
@@
  Construct $\mathbf{\Theta}$ by two unitary matrices based
       on $\mathbf{h}_{RI}$ and $\mathbf{h}_{IT}$, i.e., $\mathbf{\Theta}=\mathbf{V}_{RI}\mathbf{U}_{IT}^{\mathsf{H}}$
       with $\mathbf{V}_{RI}]_{:,1}=\frac{\mathbf{h}_{RI}^{\mathsf{H}}}{\|\mathbf{h}_{IT}\|_{2}}=\bar{\mathbf{h}}_{RI}^{\mathsf{H}}$ and
[…]
Yes $\mathcal{O}(M^{3})$
  Decouple $\mathbf{\Theta}=\mathbf{\Theta}_{0}\prod_{m=1}^{M-1}\prod_{n=m+1}^{M}\mathbf{R}_{m,n}$,
       transform to an unconstrained optimization
       and use searching-based methods,
       e.g., quasi-Newton method [de2010optimized].
No $\mathcal{O}(I_{1}M^{2}(M-1)^{2})$
  Adopt the Manifold theory to construct a
       Manifold and transform the constrained
       optimization to unconstrained optimization
       on Manifold space, on which the searching
       -based methods, e.g., Gradient-Descent,
       can be used [li2022beyond, zhao2024channel].
No $\mathcal{O}(I_{2}M^{3})$
Fully/group-connected architectures with $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$ and $\mathbf{\Theta}=\mathbf{\Theta}^{\mathsf{T}}$ Coupled unitary and symmetric constraints
  Matrix decomposition $\mathbf{\Theta}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ with
      $\mathcal{T}^{\prime}$ [nerini2023closed, sun2024new] or $\mathbf{\Theta}=\mathbf{\Psi}\mathbf{\Psi}^{\mathsf{T}}$ with $\mathcal{T}^{\prime\prime}$ [santamaria2023snr];
Yes $\mathcal{O}(M^{3})$
  Symmetric and unitary projections by
       $\mathbf{\Theta}_{\mathsf{sym}}=\frac{1}{2}(\mathbf{\Theta}_{\mathsf{opt}}+\mathbf{\Theta}_{\mathsf{opt}}^{\mathsf{T}})$ and solving
       the orthogonal Procrustes problem
       $\mathbf{\Theta}=\arg\min_{\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}}~\|\mathbf{\Theta}-\mathbf{\Theta}_{\mathsf{sym}}\|_{\mathsf{F}}^{2}$ [fang2023low];
No $\mathcal{O}(M^{3})$
  Introduce auxiliary variables $\mathbf{\Phi}=\mathbf{\Theta}$,
       with $\mathbf{\Phi}=\mathbf{\Phi}^{\mathsf{T}}$ and $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$,
       or vise versa [zhou2023optimizing].
[…]
@@
  Extract non-zero entries in $\mathbf{Y}_{I}$ and globally
       solve linear equations
[…]
Yes $\mathcal{O}(M^{3})$
  Searching-based methods, e.g., quasi-Newton
       method [shen2021] not sensitive to matrix structures.
No $\mathcal{O}\big(I_{3}\frac{M^{2}(M+1)^{2}}{4}\big)$
  Introduce auxiliary variables, e.g., $\mathbf{u}=\mathbf{h}_{RI}\mathbf{\Theta}$,
       to eliminate the matrix inverse by
       $\mathbf{\Theta}=(Y_{0}\mathbf{I}_{M}+\mathbf{Y}_{I})^{-1}(Y_{0}\mathbf{I}_{M}-\mathbf{Y}_{I})\Rightarrow$
       $\mathbf{u}(Y_{0}\mathbf{I}_{M}+\mathbf{Y}_{I})=\mathbf{h}_{RI}(Y_{0}\mathbf{I}_{M}-\mathbf{Y}_{I})$ [wu2024optimization].
[…]
@@
  • $\ddagger$
[…]
  • $\natural$
    The notation $I_{i},\forall i\in\{1,2,3\}$ denotes the number of iterations to guarantee the convergence of different methods.
  • $\natural$
[…]